OpenStudy (anonymous):

The law of cosines reduces to the Pythagorean theorem for right triangles. true or false

ganeshie8 (ganeshie8):

hint : \(\cos 90 = 0\)

ganeshie8 (ganeshie8):

law of cosines \(c^2 = a^2 + b^2 - 2ab \cos C\)

ganeshie8 (ganeshie8):



OpenStudy (anonymous):

the answer is (True )my friend ....